OpenStudy (anonymous):

Analyze the function f(x) = - 2 cot 3x. Include: - Domain and range - Period - Two Vertical Asymptotes

OpenStudy (anonymous):

@integralsabiti how would you read that can you help me read how you would saw what the domain and range and other things are

OpenStudy (anonymous):

I know the range is all real

OpenStudy (anonymous):

and I dont know what the vertical asymtote is either the two vertical asymptotes

OpenStudy (anonymous):

@amistre64 help pls :)

OpenStudy (amistre64):

since cotangent is a ratio of cos to sin; the verts are defined when sin is zero

OpenStudy (amistre64):

sin is zero at 0 and 180 degrees; so when is 3x = 0 or 180 degrees?

OpenStudy (amistre64):

the vertical asymptotes are the open circles on the number line of the link that integral provided

OpenStudy (anonymous):

im confused which one are we doing first the domain ?

OpenStudy (anonymous):

For the options @amistre64 I know the period is "pi/3" and I know The range is all real numbers I am having a hard time finding the two V.A. and the domain

OpenStudy (amistre64):

there are an infinite number of VAs, they only want you to provide 2 of them

OpenStudy (amistre64):

whenever 3x = 0 or pi, this thing becomes undefined and pops out a VA

OpenStudy (anonymous):

so what two should I say ?

OpenStudy (amistre64):

when does 3x=0? use that x value for a VA when does 3x=pi? use that x value for a VA

OpenStudy (anonymous):

can i say one is 0 and another can be 2pi n/3?? does that work?

OpenStudy (amistre64):

0 is good the "n" doesnt make it specific; 2pi/3 is fine tho

OpenStudy (amistre64):

the domain excludes all the VAs, so that would need to be generalized as: n pi/3

OpenStudy (anonymous):

The range of this equation is all real numbers and the domain is n pi/3 The Period is pi/3 and lastly the two vertical asymptotes are 0 and 2pi/3 Is this a good answer?

OpenStudy (amistre64):

domain needs some attention the way you have stated it; it says that the usable values of x that we can use are only integer values of n pi/3; when these values are actually the ones we need to avoid

OpenStudy (amistre64):

The domain is all Real values such that for any integer "n", x DOES NOT equal: n pi/3

OpenStudy (anonymous):

sorry Im confused again so I should say its all real values as well as the range or ?

OpenStudy (amistre64):

Does "domain" define all the values that we can use?

OpenStudy (anonymous):

the domain is all real numbers except vertical asymptotes (sorry for interrupting :)

OpenStudy (amistre64):

\[cot(x)=\frac{cos(3x)}{sin(3x)}~:~such~that~x\ne\frac{\pi}{3}n~for~any~integer~"n"\]why? becasue: sin(3n pi/3) = sin(n pi) = 0

OpenStudy (anonymous):

so like integersabiti said It would be all real numbers except for the VA's right @amistre64

OpenStudy (amistre64):

correct.

OpenStudy (amistre64):

except for ALL the VAs, not just the 2 that you defined

OpenStudy (anonymous):

integralsabiti means 'intagral constant ' .What you said is (integersabiti) much more interesting :D 'integer constant '

OpenStudy (anonymous):

THANK YOU SO MUCH !